If |x|+|x-1|=3 then x=?
Answer by ikleyn(53750)

Answer. This equation has two solutions: x = 2 and x = -1.

Solution

1.  If x < 0, then |x| = -x and |x-1| = -(x-1), therefore |x| + |x-1| = -x -(x-1) = -2x + 1. 

    Hence,  the equation |x| + |x-1| = 3 takes the form -2x + 1 = 3.

    The last equation has the solution x = -1,  which satisfies the inequality x < 0.

2.  If 0 <= x < 1, then |x| = x and |x-1| = -(x-1), therefore |x| + |x-1| = x -(x-1) = 1. 

    Hence,  the equation |x| + |x-1| = 3 takes the form 1 = 3.

    The last equation has no solution.

3.  If x >= 1, then |x| = x and |x-1| = (x-1), therefore |x| + |x-1| = x + (x-1) = 2x - 1. 

    Hence,  the equation |x| + |x-1| = 3 takes the form 2x - 1 = 3.

    The last equation has the solution x = 2,  which satisfies the inequality x >= 1.


The plot of the function y = |x| + |x-1| is presented in the Figure below.

      Figure. Plot y = |x| + |x-1|